The nominal interest rate is 12%. The inflation rate is 3%. The real interest rate is equal to %.
Answer #1

The formula for real interest rate is =[1+nominal interest rate/1+inflation rate] - 1

So, according to the question here, nominal interest rate is 12% and inflation rate is 3%.

Hence, [1+12/1+3]-1=[13/4]-1=9/4=2.25%

So, the real interest rate in this case is 2.25%.